# ResiDual Transformer Alignment with Spectral Decomposition

### v1

@@ A.2 Connecting TextSpan and Matching Pursuit
The second modification is that in TextSpan, the row variance of ${\bm{D}}{\bm{R}}^{T}$ is used instead of the $\ell_{1}$ norm as a criterion for atom selection. Here, we show that if this criterion is applied in SOMP (or, vice versa, the $\ell_{1}$ in TextSpan), the two algorithms are equivalent.
@@
The two dictionaries will be different: ${\bm{D}}_{MP}$ never gets updated, while ${\bm{D}}^{n}_{TS}\perp{\bm{D}}^{0}_{TS}[{\mathbb{C}}^{n}_{TS}]$ because TextSpan applies a Gram–Schmidt process that finds an orthogonal basis for the subspace of selected atoms. Since in TextSpan the dictionary is orthogonalized at each step, at time $n$ it is orthogonal to all previously chosen atoms (which are also orthogonal to each other). The dictionary of SOMP can be decomposed in two terms, one contained in the span of atoms chosen until this point by TextSpan and one orthogonal, which corresponds to the current dictionary of TextSpan:
@@
The residual of TextSpan (which, by inductive hypothesis is identical to the residual of SOMP) is by definition orthogonal to the atoms already chosen by TextSpan. Hence the selection step of SOMP will compute:
@@
Then, at step $n+1$ the two algorithms will pick the same atom index again, and update identically the residual, removing its projection on the chosen atom, i.e. ${\mathbb{C}}^{n+1}_{MP}={\mathbb{C}}^{n+1}_{TS}$. The last thing to prove is that the residual is also the same. For SOMP, the least squares solution of the optimization problem results in:
@@
A.3 Additional results
@@
A.3.1 Dataset Comparison
@@
A.3.2 Coarse Unit Alignment
CLIP-L
Dataset U U|T S R H B O
CIFAR10 0.88 0.95 0.95 0.51 0.92 0.96 0.97
CIFAR100 0.72 0.73 0.72 0.32 0.68 0.75 0.80
Cars 0.74 0.74 0.73 0.29 0.70 0.78 0.79
DTD 0.50 0.50 0.51 0.26 0.51 0.55 0.61
EuroSAT 0.71 0.71 0.73 0.36 0.53 0.62 0.95
GTSRB 0.49 0.48 0.50 0.18 0.31 0.50 0.71
MNIST 0.78 0.77 0.85 0.36 0.75 0.76 0.96
RESISC45 0.63 0.63 0.64 0.30 0.59 0.71 0.84
SUN397 0.56 0.56 0.58 0.23 0.49 0.67 0.71
SVHN 0.65 0.65 0.65 0.30 0.60 0.58 0.71
Average 0.66 0.67 0.69 0.31 0.61 0.69 0.80
@@ A.3.2 Coarse Unit Alignment
A.3.3 Spectral ResiDual Alignment

### v2

@@ A.2 Connecting TextSpan and Matching Pursuit
The second modification is that in TextSpan, the row variance of ${\bm{D}}{\bm{R}}^{T}$ is used instead of the $\ell_{1}$ norm as a criterion for atom selection. Here, we show that the two algorithms are equivalent if this criterion is applied in SOMP (or vice versa, the $\ell_{1}$ in TextSpan).
@@
The two dictionaries will be different: ${\bm{D}}_{MP}$ never gets updated, while ${\bm{D}}^{n}_{TS}\perp{\bm{D}}^{0}_{TS}[{\mathbb{C}}^{n}_{TS}]$ because TextSpan applies a Gram–Schmidt process that finds an orthogonal basis for the subspace of selected atoms. Since in TextSpan the dictionary is orthogonalized at each step, at time $n$ it is orthogonal to all previously chosen atoms (which are also orthogonal to each other). The dictionary of SOMP can be decomposed into two terms, one contained in the span of atoms chosen until this point by TextSpan and one orthogonal, which corresponds to the current dictionary of TextSpan:
@@
The residual of TextSpan (which, by inductive hypothesis, is identical to the residual of SOMP) is, by definition, orthogonal to the atoms already chosen by TextSpan. Hence the selection step of SOMP will compute:
@@
Then, at step $n+1$, the two algorithms will pick the same atom index again and update identically the residual, removing its projection on the chosen atom, i.e., ${\mathbb{C}}^{n+1}_{MP}={\mathbb{C}}^{n+1}_{TS}$. The last thing to prove is that the residual is also the same. For SOMP, the least squares solution of the optimization problem results in:
@@
A.3 Prototypical Alignment
In this section, we provide results in a different experimental setting, designed for applying our framework to unimodal encoders such as DINOv2 and ViT.
For these encoders, instead of deriving task boundaries from text encodings, we use a prototypical approach (Snell et al., 2017): for each dataset, we average the latent representations of all the corresponding samples in the training set for each target class. This produces representations that contain the average and high-level information for that class, similar to a textual encoding of a simple caption, “An image of <class>”.
DINOv2-L ViT-L
Dataset U U|T S R H B O U U|T S R H B O
CIFAR10 0.85 0.92 0.95 0.51 0.95 0.96 0.99 0.93 0.93 0.91 0.62 0.94 0.96 0.97
CIFAR100 0.77 0.79 0.78 0.25 0.84 0.88 0.89 0.68 0.70 0.70 0.34 0.81 0.81 0.86
Cars 0.43 0.43 0.44 0.05 0.46 0.43 0.72 0.24 0.26 0.27 0.07 0.36 0.28 0.48
DTD 0.65 0.65 0.65 0.30 0.67 0.70 0.76 0.50 0.50 0.51 0.28 0.59 0.58 0.67
EuroSAT 0.71 0.74 0.71 0.36 0.75 0.71 0.95 0.65 0.64 0.71 0.54 0.73 0.68 0.94
GTSRB 0.21 0.21 0.36 0.10 0.27 0.29 0.66 0.13 0.16 0.34 0.14 0.28 0.20 0.69
MNIST 0.62 0.62 0.55 0.24 0.68 0.65 0.93 0.72 0.71 0.77 0.51 0.76 0.75 0.94
RESISC45 0.76 0.75 0.76 0.25 0.78 0.74 0.87 0.50 0.47 0.50 0.29 0.68 0.57 0.79
SUN397 0.60 0.60 0.60 0.15 0.63 0.68 0.72 0.48 0.51 0.51 0.19 0.67 0.68 0.71
SVHN 0.19 0.20 0.27 0.14 0.19 0.23 0.48 0.28 0.31 0.38 0.23 0.33 0.29 0.54
Average 0.58 0.59 0.61 0.23 0.62 0.63 0.80 0.51 0.52 0.56 0.32 0.62 0.58 0.76
Figure 7: Performance comparison between image-prototype alignment methods on zero-shot classification tasks. Base: original model. Optimized: learned unit-wise weights (as in Section 4.1). ResiDual*: restricted ResiDual method as described in Section 4.2
A.4 Additional results
@@
[…]
Figure 10: Additional datasets for the study on the residual dimensionality described in Section 3.1
(a) MLP dimensionality visualization for OpenCLIP-L on ImageNet.
L20.MLP L21.MLP Picture taken in a zoo or wildlife sanctuary Serene meadow landscape Intense motorsport race Evocative streets Nostalgic vibe A shelf Balanced composition Image captured in the Patagonian wilderness Intricate architectural carving Graceful swimming fish L22.MLP L23.MLP Photo taken in Bangkok, Thailand Image showing prairie grouse Serene wilderness A photo of an adult A photograph of a big object An image of Andorra Picture taken in Rwanda Image with a zebra Picture with airplanes Picture taken in the South African safari
(b) Specializations of MLP units across layers in OpenCLIP-L.
Figure 11: Comparison of MLP dimensionality and specialization in OpenCLIP-L for ImageNet showing higher dimensionality than head units and strong superposition of concepts. (a) Visualization of MLP dimensionality. (b) Examples of MLP unit specializations across layers.
L22.H7.PC0 (high similarity) L22.H11.PC0 (low similarity)
Serene winter wonderland Photo with faded, nostalgic colors
A photo taken in the summer Classic black and white cityscape
A photo taken in the winter Caricature of a cartoon character
Image with a dramatic thunderstorm Nature macro photography
Daytime shot Ocean sunset silhouette
L22.H8.PC0 L22.H8.PC1
vast natural landscape Sublime and serene mountain lake
Futuristic technological concept Enchanting fantasy world
Picture captured in the Egyptian pyramids Image captured in the Japanese tea gardens
An irregular shape A photo with the letter L
italic text A photo with the letter Y
Figure 12: [CLS]-patches compatibility analysis. (a): Head-wise Spectral Cosine Similarity (Equation 4) between [CLS] and patch tokens for OpenCLIP-L on ImageNet. (b): Specialization of patch tokens on selected heads using OMP applied to specific principal components. L22H8, identified as a “letter” head (Figure 3), exhibits the lowest similarity in the final four layers. Interestingly, pattern-like descriptions emerge in the first principal component (PC0), while letter descriptions become apparent only in the second component (PC1).
@@
Method Atom 0 Atom 1 Atom 2
GTSRB U An acute triangle The number thirty Image with a single road sign
U|T An acute triangle The number thirty Image with a single road sign
S An acute triangle The number thirty A concentric circle
H An acute triangle The number thirty Image with a single road sign
O An acute triangle An image of the number 0 The number thirty
B An acute triangle The number thirty Serene beach sunset
MNIST U Image with three people Image with a six people A pendulum
U|T Image with three people Image with a seven people Image with a pair of subjects
S An image of three subjects Image with six subjects An image of two subjects
H An image of the number 3 A pendulum Image with a seven people
O An image of three subjects Aerial view of a coral reef detailed macro shot
B Image with a whirlpool of brimstone An image of the number 3 Playful siblings
Table 6: First 3 atoms (decompositions) obtained by SOMP on OpenCLIP-L output encodings after the different coarse unit selection methods are applied.
CLIP-L
Dataset U U|T S R H B O
CIFAR10 0.88 0.95 0.95 0.51 0.92 0.96 0.97
CIFAR100 0.72 0.73 0.72 0.32 0.68 0.75 0.80
Cars 0.74 0.74 0.73 0.29 0.70 0.78 0.79
DTD 0.50 0.50 0.51 0.26 0.51 0.55 0.61
EuroSAT 0.71 0.71 0.73 0.36 0.53 0.62 0.95
GTSRB 0.49 0.48 0.50 0.18 0.31 0.50 0.71
MNIST 0.78 0.77 0.85 0.36 0.75 0.76 0.96
RESISC45 0.63 0.63 0.64 0.30 0.59 0.71 0.84
SUN397 0.56 0.56 0.58 0.23 0.49 0.67 0.71
SVHN 0.65 0.65 0.65 0.30 0.60 0.58 0.71
Average 0.66 0.67 0.69 0.31 0.61 0.69 0.80
Figure 20: Jaccard similarity between the heads selected by the Task-conditioned Unsupervised method at 5% (introduced in Section 4.1) for each dataset on OpenCLIP-L. It shows meaningful choices, implying that for similar data distributions, similar heads are selected, and their specialization is preserved across datasets (e.g., SVHN, GTSRB, and MNIST have relevant similarities, but also EuroSAT and GTSRB).
@@ B Image with a whirlpool of brimstone An image of the number 3 Playful siblings
OpenCLIP-B OpenCLIP-L
Dataset ResiDual ResiDual ResiDual ResiDual
CIFAR10 0.97 0.97 0.98 0.98
CIFAR100 0.82 0.81 0.87 0.88
DTD 0.78 0.75 0.84 0.81
EuroSAT 0.98 0.97 0.98 0.98
GTSRB 0.88 0.84 0.92 0.91
MNIST 0.99 0.99 0.99 0.99
RESISC45 0.93 0.91 0.95 0.95
Cars 0.91 0.87 0.94 0.93
SUN397 0.77 0.72 0.82 0.80
SVHN 0.81 0.78 0.86 0.84
Average 0.88 0.86 0.92 0.91
Table 8: Accuracy produced by different configurations of ResiDual using dataset-specific bases and not ImageNet ones as in Section 4.2; (ResiDual): ResiDual in the original formulation (all principal components of all residual units); (ResiDual): ResiDual limited to head units alone (no MLPs), and PCs truncated to 90% of explained variance.